Make the variable in the bracket the subject of the given formula:
1. px-q=rx+s {x}
2. s = a/(1-r) {r}
3. A = 1/2(a+b)h {a}
1
Expert's answer
2012-09-07T11:09:03-0400
1. px-q=rx+s {x}
px-rx=s+q
x(p-r)=s+q
x=(s+q)/(p-r)

2. s = a/(1-r) {r}
s(1-r)=a
s-sr=a
sr=s-a
r=(s-a)/s

3. A = 1/2{(a+b)h} {a}
2A=ah+bh
ah=2A-bh
a=(2A-bh)/h